### v2

# Mixed-Strategy Nash Equilibrium for Crowd Navigation

Abstract
We address the problem of finding mixed-strategy Nash equilibrium for crowd navigation. Mixed-strategy Nash equilibrium provides a rigorous model for the robot to anticipate uncertain yet cooperative human behavior in crowds, but the computation cost is often too high for scalable and real-time decision-making. Here we prove that a simple iterative Bayesian updating scheme converges to the Nash equilibrium of a mixed-strategy social navigation game. Furthermore, we propose a data-driven framework to construct the game by initializing agent strategies as Gaussian processes learned from human datasets. Based on the proposed mixed-strategy Nash equilibrium model, we develop a sampling-based crowd navigation framework that can be integrated into existing navigation methods and runs in real-time on a laptop CPU. We evaluate our framework in both simulated environments and real-world human datasets in unstructured environments. Our framework consistently outperforms both non-learning and learning-based methods on both safety and navigation efficiency and reaches human-level crowd navigation performance on top of a meta-planner.
I Introduction
Figure 1: Comparison of optimality criterion for decision-making in different navigation frameworks. (a) In traditional navigation, the robot makes optimal decisions, such as to minimize the risk of collision, in a given and non-interactive environment; (b) Cooperative navigation finds optimal cooperative decisions for both the robot and the human. For decisions made based on pure-strategy Nash equilibrium, the robot expects deterministic actions from humans, which is too assertive considering the uncertain nature of human behavior; (c) Our cooperative navigation framework uses mixed-strategy Nash equilibrium as the optimality criterion, which finds probabilities of actions that represent the optimal cooperation strategies between the robot and human. The probabilistic model of mixed-strategy maintains uncertainty during interaction.
The ability to navigate fluently and safely in human-populated spaces is becoming increasingly crucial for deploying robot autonomy in real-world environments, examples include autonomous driving in populated areas [1], service robots at mass events [2], and industrial robots operating alongside human workers [3]. This problem of efficient robot navigation in human crowds with minimal disruption to humans is often referred to as crowd (social) navigation.
Human behavior studies [4, 5] reveal that humans anticipate each other’s collision avoidance behavior when navigating in crowds. Empirical results further indicate that navigation algorithms that predict human cooperation during planning could improve both safety and navigation efficiency [6]. On the other hand, coupling the prediction of cooperative human behavior with motion planning is different from multi-agent planning, such as through crowd motion simulation [7, 8], where humans follow pre-defined action rules. Instead, both the humans and the robot should be modeled as optimal planners, whose individual objectives depend on others‘ actions [9]. As a result, the simultaneous prediction of cooperative human behavior and planning of robot motion become the outcome of assessing the coupled optimal actions of all the agents. This model leads to the application of game theory models in crowd navigation, specifically the application of Nash equilibrium [10] as an optimality criterion for decision-making.
In a Nash equilibrium, no agent wants to change their strategy unless others do the same, since each agent’s strategy is optimal given other agents’ Nash equilibrium strategies. There are two kinds of strategies that can create a Nash equilibrium, pure-strategy and mixed-strategy, and we are interested in the latter. Pure-strategy refers to choosing deterministic and specific actions, while mixed-strategy refers to randomly sampling actions from a distribution of pure-strategies. Crucially, John Nash proved that not all games have a pure-strategy Nash equilibrium, but at least one mixed-strategy Nash equilibrium exists for all games [11]. Our interest in mixed-strategy Nash equilibrium stems from not only its mathematical rigor but also its non-deterministic nature. Prediction based on pure-strategy Nash equilibrium would expect humans to follow the paths exactly as predicted, which is too assertive given the uncertain nature of human behavior. Mixed-strategy Nash equilibrium, however, maintains such behavioral uncertainty during the robot’s decision-making. A high-level comparison of traditional navigation and cooperative navigation with pure-strategy Nash equilibrium and mixed-strategy Nash equilibrium is shown in Figure 1.
Despite its potential, applying mixed-strategy Nash equilibrium to crowd navigation faces several challenges. First, computing mixed-strategy Nash equilibrium is often considered impractical due to its computation burden—even for a 3-agent game, the computation can be close to NP-complete [12]. Second, existing Nash equilibrium frameworks often focus on discrete actions and a finite number of strategies (e.g., poker). However, crowd navigation studies physical agents navigating in a continuous space, with extra physical constraints. Third, the practical implication of Nash equilibrium heavily depends on how well the game design matches human behavior, yet there is no consensus for a proper game model for crowd navigation.
In this work, we formally prove that—by modeling an agent’s mixed-strategy as a probabilistic belief of the agent’s intended path—a simple iterative Bayesian updating scheme guarantees the convergence to the global Nash equilibrium. In the corresponding game, each agent aims to minimize the expected risk of collision with other agents, as well as the deviation from a nominal mixed-strategy. Furthermore, we propose a framework to characterize the nominal mixed-strategy of each agent by extracting a temporal Gaussian process from offline trajectory datasets. This data-driven characterization process anecdotally captures the difference in crowd navigation patterns between humans and mice. Based on these theoretical results, we propose a sampling-based crowd navigation algorithm. The algorithm has a lower time complexity with respect to the number of agents compared to the state-of-the-art dynamic game solvers and runs in real-time on a laptop CPU. Given measurements of pedestrian positions and velocities, the algorithm infers mixed-strategy Nash equilibrium for up to 8 agents as a distribution of potential cooperation collision avoidance behaviors and chooses the optimal navigation decision for the robot accordingly. The algorithm is further integrated into a complete navigation stack for crowd navigation in unstructured real-world environments with static occupancy grid maps. Figure 2 provides a high-level illustration of the proposed iterative Bayesian updating scheme for computing mixed-strategy Nash equilibrium in crowd navigation.
Our work diverges from existing game-theoretic crowd navigation methods. Existing methods follow a top-down approach, where no assumption about game objective structure and game decision space is made in advance—these components will be specified later for applications like crowd navigation. While the solution methodologies would apply to any game, not just crowd navigation game, this generality comes at the price of local optimal solvers [^1] Global Nash equilibrium solvers for arbitrary games is computationally intractable. and unaffordable computation cost and complexity for real-time decision-making in human crowds. In contrast, we take a bottom-up approach, where we formulate a specific game based on unique behavioral features in crowd navigation, such as goal-oriented and collision avoidance-driven behavior. This bottom-up approach provides a sufficient structure to develop a planner with game theory guarantees, including recovering global mixed-strategy Nash equilibrium, without compromising the computation efficiency for real-time decision-making. As demonstrated in our experiments, the computation efficiency of our method enables rapid replanning of the robot’s paths, which plays a critical role in improving navigation performance when interacting with potentially non-cooperative pedestrians.
The rest of the paper is organized as follows: Section II reviews the literature on crowd navigation and discusses the key difference between our method and existing methods. We first introduce the iterative Bayesian updating scheme in Section III. Then we show how this iterative updating scheme is guaranteed to converge to a Nash equilibrium of a mixed-strategy crowd navigation game in Section IV, with extra safety property analysis of the Nash equilibrium. Section V introduces the method to learn nominal strategies from trajectory datasets. Section VI discusses how to use the mixed-strategy Nash equilibrium to infer human anticipation in practice with sampling-based approaches. Section VII contains the details of the evaluation results. Lastly, we conclude the paper and provide further discussion in Section VIII.
II Related work
Figure 2: A high-level illustration of our framework. From left to right: The perception stack first estimates the states of the agents (e.g. robot and human positions). Then with the temporal Gaussian process kernel learned from offline datasets, we construct the initial mixed-strategies for the agents as Gaussian processes. After that, we apply the iterative Bayesian updating scheme to compute the mixed-strategy Nash equilibrium.
II-A Early work on robot navigation in crowds
Roboticists have been investigating navigation in human environments since the 1990s. Two landmark studies were the RHINO [13] and MINERVA [14] experiments, where robotics systems were deployed in museums to provide tour-guide to thousands of visitors. Additional work for tour-guide robots was also conducted later, such as Robox [15], Mobot [16], Rackham [17], and CiceRobot [18]. A comprehensive review of the history of crowd navigation can be found in [19].
These works use conventional indoor navigation stacks such as the dynamic window approach [20], where humans are modeled as non-reactive obstacles. While these methods are sufficient when the robot interacts with sparse crowds, the robot’s navigation efficiency is limited because of safety concerns. For example, the robot’s speed is intentionally limited to avoid constant emergency stopping [16].
II-B Prediction-then-planning methods
The limitations of conventional navigation methods in human crowds motivate researchers to develop human-aware navigation methods. One prevalent framework in this category is to predict human motion and plan robot actions to avoid hindering human motion, we name such framework as prediction-then-planning. Human motion can be predicted in the form of cost maps, such as from inverse reinforcement learning [21, 22] or unsupervised learning [23]. Human motion can also be directly predicted in forms of trajectories with deep learning methods, such as using long-short term memory models (LSTM) [24], generative adversarial networks (GAN) [25] or graph-based models [26]. We refer readers to [27] for a comprehensive review of human trajectory prediction. Note that, [28] reports that a simple constant velocity model could outperform deep learning-based trajectory prediction methods, arguing that existing neural network architectures are insufficient at capturing interpersonal interaction. In addition, another thread of research aims to develop fast reactive planners around dynamic obstacles, such as through field representations [29] or graph search [30].
Planning on top of human motion prediction also introduces the necessity of uncertainty-aware planning: motion planning needs to take into account predictive uncertainty in human motion. In [31], closed-loop belief updating of dynamic obstacles’ states is incorporated into receding horizon planning, which reduces the anticipated obstacle uncertainty and generates less conservative navigation behavior. A confidence-ware motion planning framework is proposed in [32] that maintains a Bayesian belief of the human motion prediction confidence with online observations. By explicitly modeling the confidence of prediction, the planner can be robust against unexpected human behavior. Robust motion planning with uncertain human motion prediction has also been investigated in other works. In particular, [33] introduces a decoupled framework that combines learning-based trajectory prediction [26] with risk-sensitive model predictive control [34]. Lastly, a flipped approach is introduced in [35] in the context of autonomous driving, where a risk-sensitive trajectory prediction approach is proposed for robust planning.
Limitations of prediction-then-planning methods
While demonstrating more compliant navigation behavior alongside humans, decoupled prediction and planning is limited by its core assumption—that the robot’s action will not influence humans’ actions. Failing to account for human reaction could lead to robot actions that surprise humans, who in turn react out of the robot’s expectation, resulting in short oscillatory interaction, often referred to as “reciprocal dance” [36]. Failing to account for human reaction in uncertainty quantification could also lead to over-conservative navigation robot behavior—without incorporating human reaction to belief updating during planning, the over-approximated predictive uncertainty could lead the robot to consider all viable paths are unsafe and the only safe option is to stay still, a phenomenon often referred as the “freezing robot problem” [6].
II-C Coupled prediction and planning
Given the limitations of decoupled prediction and planning, it becomes necessary to lift the assumption that the robot does not interfere with human intents. Thus, an alternative framework, named coupled prediction and planning, seeks to simultaneously plan robot actions and predict human actions. The seminal work [37] shows that cooperative collision avoidance (CCA)—where humans and robots collectively make decisions for collision avoidance—is crucial for avoiding artifacts such as “reciprocal dance” and “freezing robot problem”. The importance of CCA in dense crowds is experimentally verified in a psychophysical study [4].
In [6], an individual’s intent is modeled as a Gaussian process and CCA is modeled as a joint decision-making process by coupling Gaussian processes through a collision avoidance-based likelihood function. The statistical optimality of coupled Gaussian processes is further investigated in [38, 39]. In [40], a distribution space crowd navigation model is proposed, with agent intent modeled as a distribution of trajectories. While similar to the mixed-strategy model in this work, the model in [40] is not a game-theory model but is instead a joint decision-making model similar to [6]. Topology-based abstractions are also used for modeling CCA, such as through braids [41, 42] or topological-invariance [43, 44, 45]. CCA is also studied from the perspective of opinion dynamics [46], with an explicit focus on breaking deadlock situations such as “reciprocal dance” [47].
Another group of works focuses on the implicit modeling of cooperative collision avoidance. Implicit CCA models are often obtained from data from real-world or simulated human pedestrians—the availability of data will impact the choice of modeling technique. Given the cost of human data collection, implicit modeling with human data cannot afford directly training robot navigation policies with humans, instead uses techniques such as imitation learning [48] or inverse reinforcement learning [49, 50]. Instead of collecting human data, an alternative solution is to simulate cooperative navigation agents using decentralized multi-agent collision avoidance methods, such as [7, 8]. The benefit of implicit CCA modeling from simulated data is that the navigation policies can be trained directly with the reactive agents using reinforcement learning methods, such as [51, 52, 53, 54]. However, the sim-to-real transfer of such policies remains an open challenge, mainly due to the distribution shifts between simulated crowd behavior and real-world human behavior. It is also unclear whether the reinforcement learning policies explicitly predict human cooperation during planning, instead of treating humans merely as dynamic obstacles—the latter falls into the category of the prediction-then-planning framework.
Note that there is a relevant group of works on social convention-aware navigation. The goal is to take into account social conventions during motion planning, such as people tend to walk in groups [55, 56] or people would prefer not to move when forming certain geometrical structures [57]. Though conceptually similar, we consider this body of work to be essentially different from coupled prediction and planning, because the social conventions are defined a priori. As a result, these planners do not necessarily take into account human reaction to the robot’s actions.
Limitation of existing coupled-prediction-and-planning methods
The main limitation of existing coupled-prediction-and-planning methods is the lack of a rigorous cooperative model, given humans and the robot are cooperating through only mutual observation instead of explicit communication. Cooperation is often granted in coupled-prediction-and-planning methods. For example, methods that formulate crowd navigation as a joint decision-making problem assume that humans and robots share the same joint decision-making objective. We argue that each agent in crowd navigation should be modeled as an independent decision-making individual, the phenomenon of cooperative collision avoidance is an emergent behavior from each agent optimizing their individual objective, which partially depends on other agents’ actions. This insight naturally leads to the modeling of crowd navigation through game theory.
II-D Game-theoretic planning
Even though game-theoretic planning falls into the category of coupled prediction and planning, here we separately discuss this body of work given its close connection to our work. The key difference between game-theoretic methods and other coupled-prediction-and-planning methods is that game theory assumes each agent makes individual optimal decisions and it provides a rigorous optimality criterion for decision-making—the concept of equilibrium. The two most commonly used equilibrium notions are Nash equilibrium [10] and Stackelberg equilibrium [58]. Stackelberg equilibrium, in general, has lower computation complexity since it enforces a leader-follower structure. On the other hand, Nash equilibrium is considered as the more rigorous and natural notion, since it assumes equal status for all agents and all agents make decisions simultaneously.
In order to apply game theory models, crowd navigation is often formulated as a dynamic game: agents, with states governed by certain dynamics, optimize control policies as pure-strategies. One of the pioneering works in this direction is [9], where the interaction between a human-driven vehicle and an autonomous vehicle is modeled as a dynamic game. Since then, dynamic game solvers with better numerical efficiency have been developed. [59] formulates a general-sum dynamic game for crowd navigation, where the local Nash equilibrium is solved through the combination of an iterative best response scheme and iterative linear-quadratic regulator. [60] proposes a fast augmented Lagrangian solver that converges locally to the generalized Nash equilibrium of a dynamic game, which supports real-time model predictive control for autonomous driving. However, pure-strategy Nash equilibrium expects humans to react exactly as predicted, which is too assertive for interacting with humans. Furthermore, existing pure-strategy game solvers can only approximate a locally optimal solution, since computing the Nash equilibrium of a generalized game is computationally intractable [12].
To address the lack of flexibility in pure-strategy Nash equilibrium, probabilistic variants of Nash equilibrium have been proposed. As a pioneer work in this direction, [61] proposes a stochastic dynamic game formula that lifts the strict rationality assumption of Nash equilibrium and instead assumes bounded rationality of agents. The authors propose a new notion of equilibrium named Entropic Cost Equilibrium (ECE) and show that ECE is equivalent to the mixed-strategy Nash equilibrium of a maximum entropy game. The proposed equilibrium formula also enables inverse inference of interaction policy from observations. In [62, 63], the multi-modality in stochastic games is investigated from the perspective of linear-quadratic games and partially observable Markov decision process (POMDP). In [64], an explicit mixed-strategy game formula is proposed, allowing agents to simultaneously optimize for multi-agent pure-strategies, but this formula does not recover the mixed-strategy Nash equilibrium.
Lastly, game theory models can be combined with other methods for online adaptation during crowd navigation. For example, Nash equilibrium can be used to infer other agents’ internal states, such as altruism. [65] proposes to use Nash equilibrium as an inference model for online estimation of human drivers’ social value orientation, leading to the improved prediction accuracy of human driver actions during the interaction. Similar ideas are also explored in [66, 67, 3]. Another example of bootstrapping adaptive decision-making with game-theoretic models is [68], where the game-theoretic model is used for providing candidates for contingency planning.
Limitations of existing game-theoretic planning methods
There are two major limitations of existing game-theoretic planning methods for crowd navigation. The first limitation is the high computation complexity, which prevents existing methods from being applied to real-time navigation in human crowds. For example, as one of the fastest pure-strategy dynamic game solvers, ALGAMES [60] still suffers from a time complexity of $O(M^{3})$ with $M$ being the number of players, thus can only perform real-time inference with no more than 2 human agents. The computation of mixed-strategy Nash equilibrium is even more burdensome. In [61], the method is also only demonstrated with no more than 2 human agents. In [64], the computation of mixed-strategies is offloaded to a neural network training offline and this approach is only demonstrated for two-agent interaction. The second limitation is the lack of flexibility in representing uncertainty in strategies. In [61] is based on the Linear-Quadratic-Gaussian formula. In [64], the mixed-strategy model is limited to a small finite number of pure-strategies, instead of a probability distribution.
We argue that these two limitations are largely due to the “top-down” framework taken by existing methods. In the “top-down” framework, multi-agent interaction is formulated and solved as a generalized game, where no assumption of agent objective is made. The agent objective will be specified later for specific applications such as crowd navigation. While this “top-down” approach could be applied for arbitrary types of multi-agent interaction beyond crowd navigation, it suffers from high computation complexity and often requires narrowing down strategy representation (e.g., Gaussian distributions). Our work, however, takes a “bottom-up” approach—we tailor a stochastic game formula specifically for crowd navigation, based on unique properties in crowd navigation, such as the assumption that each agent has an individual goal-oriented objective while sharing a collision avoidance objective with other agents. This “bottom-up” approach allows us to apply calculus of variations [69] to compute mixed-strategy Nash equilibrium in closed form, which further enables real-time inference with a larger number of agents (up to 8 agents on a laptop CPU). In addition, our algorithm and proofs support arbitrary types of distribution and collision avoidance cost functions, which enriches the representation flexibility of our model.
III Bayes’ rule for crowd navigation
In this section, we introduce a simple Bayesian updating scheme for crowd navigation.
III-A Notations and definitions
We assume there are $N$ agents, including the robot and the pedestrians, in a two-dimensional navigation environment. We start by defining a set of unique indices =$\mathcal{I}=\{1,2,\dots,N\}$ for all agents, where index $1$ is reserved for the robot. The state space of each agent is denoted as $\mathcal{X}\subset\mathcal{R}^{2}$, as we are primarily interested in agents’ planar positions. We model agent $i$’s pure-strategy as a continuous trajectory $s_{i}:\mathbb{R}^{+}\mapsto\mathcal{X}$, which maps time to the agent’s state at that moment. The space of all possible pure-strategies is called trajectory space, denoted as $\mathcal{S}$.
Remark 1.
We do not address the perception aspect of the crowd navigation problem. In this work, we assume the robot has reasonably accurate estimations of its pose and nearby pedestrians’ positions.
Definition 1.
(Mixed-strategy) We model a mixed-strategy as a probabilistic belief over an agent’s possible pure-strategies. We denote agent $i$’s mixed-strategy as:
$p_{i}(s):\mathcal{S}\mapsto\mathbb{R}_{0}^{+}.$ (1)
Lemma 1.
Since a mixed-strategy is a probability distribution, it must satisfy the following property:
[…]
Definition 2.
(Nominal mixed-strategy) We define each agent has a nominal mixed-strategy before interacting with other agents, denoted as $p^{\prime}_{i}(s)$.
Remark 2.
We assume the nominal mixed-strategy represents the agent’s intent without the presence of other agents, thus it does not reflect collision avoidance behavior.
Definition 3.
(Collision risk) Collision risk is defined as a function $r(s_{1},s_{2}):\mathcal{S}\times\mathcal{S}\mapsto\mathbb{R}_{0}^{+}$, which evaluates the collision risk between two pure-strategies (trajectories).
Assumption 1.
We assume the collision risk function’s support is bounded, close, and compact.
Definition 4.
(Expected collision risk) Given one agent’s mixed-strategy $p_{i}$, the expected risk of another given pure-strategy $s$ colliding with the agent is:
[…]
Definition 5.
(Joint expected collision risk) The joint expected collision risk between two agents is defined as the joint expectation of collision risk with respect to their mixed-strategies:
[…]
Figure 3: Examples of the negotiation convergence process during our iterative Bayesian updating scheme. (a) Two-agent negotiation in one-dimensional space. (b) Two-agent hallway passing. Here the mixed-strategy is modeled in the space of trajectories. (c) Four-agent crossing.
III-B Two-agent social navigation
We start with describing the iterative Bayesian updating scheme that finds the mixed-strategy Nash equilibrium between two social navigation agents. Bayesian updating requires two key components: a prior belief and a conditional likelihood. Here the prior belief is each agent’s current mixed-strategy, denoted as $p_{i}^{\prime}(s)$. By assuming each agent follows the expected utility hypothesis [70], we propose the following conditional likelihood function to reflect collision avoidance behavior.
Definition 6.
(Conditional likelihood) Given the mixed-strategy of agent $j$, the likelihood of measuring agent $i$ following a pure-strategy $s$ is defined as:
$z(s|p_{i})$ $=\exp\Big{(}{-}\mathbb{E}_{p_{j}}[r](s)\Big{)}.$ (5)
Remark 3.
The use of inverse exponential in (5) is an empirical design choice, inspired by Gaussian distributions. But as will be shown later, the inverse exponential is necessary for the convergence to Nash equilibrium.
Definition 7.
(Bayesian posterior strategy) Given the mixed-strategy $p_{i}^{\prime}(s)$ of agent $i$ and the mixed-strategy $p_{j}(s)$ of agent $j$, the Bayesian posterior mixed-strategy of agent $i$ after applying Bayes’ rule is:
$p_{i}(s)$ $=\eta\cdot p_{i}^{\prime}(s)\cdot z(s|p_{j})$ (6)
$=\eta\cdot p_{i}^{\prime}(s)\cdot\exp\Big{(}{-}\mathbb{E}_{p_{j}}[r](s)\Big{)}$ (7)
where $\eta$ is the normalization term.
We now describe the iterative Bayesian updating algorithm for two-agent social navigation. Figure 3(a) and Figure 3(b) show iterations of two-agents’ mixed-strategies in an illustrative one-dimensional space and a hallway passing scenario, respectively.
Algorithm 1 Two-agent social navigation
1:procedure TwoAgentSocialNav($p_{i}^{\prime},p_{j}^{\prime}$)
2:     $k\leftarrow 0$triangleright $k$ is the negotiation step.
3:     $p_{i}^{[k]}\leftarrow p_{i}^{\prime}$
4:     $p_{j}^{[k]}\leftarrow p_{j}^{\prime}$
5:     while convergence criterion not met do
6:         $p_{i}^{[k+1]}(s)\leftarrow\eta\cdot p_{i}^{\prime}(s)\cdot z(s|p_{j}^{[k]})$
7:         $p_{j}^{[k+1]}(s)\leftarrow\eta\cdot p_{j}^{\prime}(s)\cdot z(s|p_{i}^{[k+1]})$
8:         $k\leftarrow k+1$
9:     end while
10:     return $p_{i}^{[k]}$ and $p_{j}^{[k]}$
11:end procedure
Algorithm 2 Multi-agent crowd navigation
1:procedure MultiAgentSocialNav($p^{\prime}_{1},\dots,p^{\prime}_{N}$)
2:     $k\leftarrow 0$triangleright $k$ is the negotiation step.
3:     for $i\in[1,N]$ do
4:         $p_{i}^{[k]}(s)\leftarrow p^{\prime}_{i}(s)$
5:     end for
6:     while convergence criterion not met do
7:         for $i\in[1,N]$ do
8:              $p^{[k]}_{/i}\leftarrow\frac{1}{N-1}\left[{\sum_{j<i}}p^{[k+1]}_{j}{+}\sum_{j>i}p^{[k]}_{j}\right]$
9:              $p_{i}^{[k+1]}\leftarrow\eta\cdot p^{\prime}_{i}\cdot z\left(s|p^{[k]}_{/i}\right)$
10:         end for
11:         $k\leftarrow k+1$
12:     end while
13:     return $p_{1}^{[k]},\dots,p_{N}^{[k]}$
14:end procedure
III-C Multi-agent crowd navigation
We now extend Algorithm 1 to an arbitrary number $N$ of agents. We first extend the definition of the conditional likelihood function.
Definition 8.
(Conditional likelihood (multi-agent)) Given the mixed-strategies of all agents other than $i$, the likelihood of measuring agent $i$ following a pure-strategy $s$ is defined as:
$z(s|p_{/i})$ $=\exp\Big{(}{-}\mathbb{E}_{p_{/i}}[r](s)\Big{)}.$ (8)
$p_{/i}(s)$ $=\overline{\sum_{j\in\mathcal{I}/i}}p_{j}(s)$ (9)
where $\overline{\sum}$ denotes averaged sum.
Definition 9.
(Bayesian posterior strategy (multi-agent)) Given the mixed-strategy $p_{i}^{\prime}(s)$ of agent $i$ and the mixed-strategies of all the rest of the agents, the Bayesian posterior mixed-strategy of agent $i$ after applying Bayes’ rule is:
$p_{i}(s)$ $=\eta\cdot p_{i}^{\prime}(s)\cdot z(s|p_{/i})$ (10)
$=\eta\cdot p_{i}^{\prime}(s)\cdot\exp\Big{(}{-}\mathbb{E}_{p_{/i}}[r](s)\Big{)}$ (11)
where $\eta$ is the normalization term.
We now describe the iterative Bayesian updating algorithm for multi-agent crowd navigation in Algorithm 2. Figure 3(c) shows iterations of multi-agent mixed-strategies in a crossing scenario.
Across all the examples in Figure 3, we want to point out two important properties of our algorithm: (1) Bayesian updating allows non-symmetric and multi-modal mixed-strategies during iterations, which expands the expressiveness of our method; (2) Although the iterative Bayesian updating scheme does not simultaneously update all agents’ strategies, the converged mixed-strategies can still be symmetrical. As we will discuss in the next section, the symmetry comes from the fact that our algorithm converges to a global Nash equilibrium. A detailed demonstration of mixed-strategy evolution during the iterative Bayesian updating scheme can be found in Extension 1.
IV Mixed-strategy Nash equilibrium for social navigation
We now show that Algorithm 2 is guaranteed to converge to a global Nash equilibrium of a general-sum mixed-strategy game. Furthermore, the converged Nash equilibrium guarantees a lower-bounded reduction of the joint expected collision risk among all agents.
Definition 10.
(General-sum game) We define a $N$-player general-sum mixed-strategy game. The objective of player $i$ is:
$J_{i}(p_{1},\cdots,p_{N})$ $=\mathbb{E}_{p_{i},p_{/i}}[r]+D(p_{i}\|p^{\prime}_{i})$ (12)
$p_{/i}(s)$ $=\overline{\sum_{j\in\mathcal{I}/i}}p_{j}(s)$ (13)
where $D(\cdot\|\cdot)$ is the Kullback-Leibler(KL)-divergence between two distributions. We assume the nominal mixed-strategies $p^{\prime}_{i}$ are given a priori and accessible by all players.
Theorem 1.
Algorithm 2 always converges. When $p^{\prime}_{i}=p_{i}^{\prime},\forall i\in\mathcal{I}$, the converged mixed-strategies $(p_{1}^{*},\dots,p_{N}^{*})$ assemble a global Nash equilibrium of the general-sum game, meaning the following equation holds for any $i\in\mathcal{I}$:
$p_{i}^{*}$ $=\operatorname*{arg\,min}_{p_{i}}J_{i}(p_{1}^{*},\dots,p_{i},\dots,p_{N}^{*})$
Proof.
We leave details of the full proof in the appendix. The proof depends on the following theorem, which by itself reveals another important property of the Bayesian updating scheme. ∎
Theorem 2.
The Bayesian posterior in two-agent social navigation (7) is the global minimum of the following optimization problem:
$\eta\cdot p_{i}^{\prime}(s)\cdot\exp$ $\Big{(}{-}\mathbb{E}_{p_{j}}[r](s)\Big{)}$
$=\operatorname*{arg\,min}_{p}\mathbb{E}_{p,p_{j}}[r]+D(p\|p_{i}^{\prime})$ (14)
[…]
Proof.
We can expand the objective function:
$\mathbb{E}_{p,p_{j}}[r]+D(p\|p^{\prime}_{i})$ (16)
[…]
This optimization problem can be solved as an isoperimetric problem with a subsidiary condition, the Lagrangian of which is:
$\mathcal{L}(p,\lambda)$
[…]
where $\lambda\in\mathbb{R}$ is the Lagrange multiplier. Based on the Lagrangian, the necessary condition for $p^{*}(s)$ to be an extreme is (Theorem 1, Page 43 [69]):
$\frac{\partial\mathcal{L}}{\partial p}(p^{*},\lambda)$ $=\log p^{*}(s)+1-\log p^{\prime}_{i}(s)+\mathbb{E}_{p_{j}}[r](s)-\lambda$
$=0$
$p^{*}(s)$ $=p^{\prime}_{i}(s)\cdot\exp\Big{(}{-}\mathbb{E}_{p_{j}}[r](s)+\lambda-1\Big{)}$
$=p^{\prime}_{i}(s)\cdot\exp\Big{(}{-}\mathbb{E}_{p_{j}}[r](s)\Big{)}\cdot\exp(\lambda-1)$ (19)
By substituting (19) to the equation constraint (15), we have:
[…]
$=1$
[…]
Substituting (21) back to (19) gives us:
[…]
$=\eta\cdot p^{\prime}_{i}(s)\cdot z(s|p_{j})$ (23)
Since the objective function (17) is unbounded from above, this extremum $p^{*}(s)$ is the global minimum of the problem. ∎∎
Theorem 2 can be extended to the Bayesian posterior of multi-agent crowd navigation (11).
Theorem 3.
The Bayesian posterior in multi-agent social navigation (11) global minimum of the following optimization problem:
$\eta\cdot p_{i}^{\prime}(s)\cdot\exp$ $\Big{(}{-}\mathbb{E}_{p_{/i}}[r](s)\Big{)}$
$=\operatorname*{arg\,min}_{p}\mathbb{E}_{p,p_{/i}}[r]+D(p\|p_{i}^{\prime})$ (24)
[…]
Proof.
Follow the same proof of Theorem 2, with the distribution $p_{j}$ being replaced by $p_{/i}$. ∎∎
Remark 4.
The objective function (17) is the summation of two sub-objectives that represents the two most important behavioral properties in crowd navigation: collision avoidance and goal-orientated behavior. The first sub-objective $\mathbb{E}_{p,p_{i}}[r]$ evaluates the expected risk of collision between two agents. On the other hand, minimizing the second objective—the KL-divergence with respect to the nominal mixed-strategy—preserves the agent’s original intent during collision avoidance.
Remark 5.
Theorem 3 indicates that Algorithm 2 is conceptually similar to the commonly used iterative best response (IBR) scheme. However, IBR does not guarantee convergence, nor does it recover the global Nash equilibrium, while our iterative scheme is guaranteed to recover a global mixed-strategy Nash equilibrium.
On top of the guaranteed convergence to the mixed-strategy Nash equilibrium, we show that the converged mixed-strategies also guarantee a lower-bounded reduction in expected collision risk.
Theorem 4.
The converged mixed-strategies $(p_{1}^{*},\dots,p_{N}^{*})$ from Algorithm 2 has a lower joint expected collision risk, compared to the nominal mixed-strategies. Furthermore, the reduction in expected collision risk is lower-bounded by the summation of KL-divergence between each agent’s converged mixed-strategy and the nominal mixed-strategy:
$\sum_{i=1}^{N}\sum_{j=i+1}^{N}\mathbb{E}_{p_{i}^{\prime},p_{j}^{\prime}}[r]-\sum_{i=1}^{N}\sum_{j=i+1}^{N}\mathbb{E}_{p_{i}^{*},p_{j}^{*}}[r]\geq\sum_{i=1}^{N}D(p_{i}^{*}\|p_{i}^{\prime})$ (26)
Proof.
See appendix. ∎
Remark 6.
Our model is highly generalizable for social navigation, since Algorithm 2, Theorem 1 and Theorem 4 do not depend on any assumption on the specific form of mixed-strategy and collision risk function, thus are compatible with arbitrary types of mixed-strategies and collision risk function. On the other hand, the game objective structure (12) is specifically for social navigation—minimizing the KL-divergence between the converged mixed-strategy from the nominal strategy does not necessarily make sense for other game-theoretic interactions.
Figure 4: Demonstration of learning temporal kernels from trajectory datasets and how they influence the Bayes’ rule Nash equilibrium. In each column of (a) and (b), we show from left to right: The snapshots from the original human [71] and mouse dataset [72]; Temporal correlation from the transformed variance dataset (45) and the optimal kernel after training; Trajectory samples from the nominal mixed-strategies characterized by the learned kernels; Trajectory samples from the Nash equilibrium mixed-strategies.
V Learning gaussian process model for nominal mixed-strategy
In Algorithm 2 we assume the nominal mixed-strategy of each agent is given. This section describes constructing mixed-strategy as a Gaussian process (GP). A Gaussian process is a Gaussian distribution over functions [73], in our case, trajectories. Similar to finite-dimensional multivariate Gaussian distributions, a GP can be characterized by a mean function $m(t)$ and a covariance kernel $K(t,t^{\prime})$:
$s(t)\sim\mathcal{GP}\big{(}m(t),K(t,t^{\prime})\big{)}$ (27)
Since we model trajectory as a mapping from time to agent state, the Gaussian processes in this work are temporal stochastic processes
Assumption 2.
Throughout the paper, we assume we have a symmetric and stationary kernel $K(t,t^{\prime})$:
[…]
We first discuss choosing the GP mean function for each agent’s nominal mixed-strategy. Here we construct the mean function for humans as a constant velocity model based on the current measured velocity [^2] As pointed out in [28], a constant velocity model could outperform state-of-the-art learning-based methods for short-term trajectory prediction. Especially since we assume nominal mixed-strategies do not encode collision behavior, the constant velocity model is a suitable model in our case.. For the robot, we use a high-level meta-planner to plan a path toward the goal as the mean of the nominal mixed-strategy. In an unstructured environment, the meta-planner can be the standard path planner (e.g., RRT) on top of the occupancy grid map of the environment.
With the GP mean functions constructed, the characterization of GPs as nominal mixed-strategies is equivalent to the characterization of GP kernel functions. Here, we introduce a method for learning GP kernel from offline trajectory datasets based on agents’ preference for maintaining a constant velocity. We extract this preference by measuring the temporal correlation across different times within a trajectory.
V-A Temporal correlation and conditional variance in GPs
We define temporal correlation as the correlation between the agent states across different times. Based on the marginal property of Gaussian processes, given a GP denoted as $\mathcal{GP}(m(t),K(t,t^{\prime}))$, we have that the marginal distribution of the GP at any two time steps $(t_{1},t_{2})$ is a finite-dimensional Gaussian distribution:
$\begin{bmatrix}s(t_{1})\\ s(t_{2})\end{bmatrix}$ $\sim\mathcal{N}\left(\begin{bmatrix}m(t_{1})\\ m(t_{2})\end{bmatrix},\begin{bmatrix}K(t_{1},t_{1})&K(t_{1},t_{2})\\ K(t_{2},t_{1})&K(t_{2},t_{2})\end{bmatrix}\right)$ (34)
$=\mathcal{N}\left(\begin{bmatrix}m_{1}\\ m_{2}\end{bmatrix},\begin{bmatrix}K_{11}&K_{12}\\ K_{21}&K_{22}\end{bmatrix}\right)$ (39)
The covariance matrix entries $K_{11},K_{22}$ determine the uncertainty of the agent state at the time $t_{1}$ and $t_{2}$, respectively. The off-diagonal entry $K_{12}=K_{21}$ determines the correlation between the agent states at time $t_{1}$ and at time $t_{2}$. Furthermore, based on the conditional property of multivariate Gaussian distribution, the Gaussian distribution of the state $x_{2}=s(t_{2})$ at time $t_{2}$, conditioned on the state $x_{1}=x(t_{1})$ at time $t_{1}$, can be represented through the off-diagonal entry $K_{12}=K_{21}$:
$x_{2}\sim\mathcal{N}\Big{(}m_{1}{+}K_{12}K_{22}^{-1}(x_{1}{-}m_{2}),K_{11}{-}K_{12}K_{22}^{-1}K_{21}\Big{)}$ (40)
We can generalize the above formula of variance of conditional Gaussian in finite-dimensional space to the infinite-dimensional space of Gaussian processes.
Definition 11.
(Conditional variance in GP) Given a temporal covariance kernel function $K(\cdot,\cdot)$, (40) is the conditional variance between two time steps $t_{1},t_{2}$:
$c(t_{1},t_{2})=K(t_{1},t_{1}){-}K(t_{1},t_{2})K(t_{2},t_{2})^{-1}K(t_{2},t_{1})$ (41)
Lemma 2.
Based on Assumption 2, the covariance kernel is stationary and symmetric, thus the conditional variance defined in (41) is stationary and symmetric:
[…]
Conditional variance (41) essentially transforms a high-dimensional probability distribution of trajectories (GP) to a low-dimensional space of displacement variance, enabling a way to extract a covariance kernel from offline datasets.
V-B Learning GP kernel from datasets
To characterize GP kernels from trajectory datasets, we can optimize the kernel parameter, such that the conditional variance of the GP as a function (41) optimally matches the conditional variance estimated from trajectory datasets.
Consider a trajectory dataset, denoted as $\mathcal{T}=\{s_{i}\}_{i\in\mathcal{I}}$, where agent $i$’s trajectory across $T$ time steps is a vector $s_{i}=[x_{i}^{[1]},\dots,x_{i}^{[T]}]\in\mathbb{R}^{T\times 2}$. Given a difference between two time steps $\delta t$, we can transform the original trajectory dataset $\mathcal{T}$ into a displacement dataset $\mathcal{D}$:
$\mathcal{D}_{\delta t}=\{\Delta x_{i}^{[t]}\},\quad\Delta x_{i}^{[t]}=|x_{i}^{[t+\delta t]}-x_{i}^{[t]}|$ (44)
The displacement dataset $\mathcal{D}_{\delta t}$ can transformed into a variance dataset $\mathcal{V}_{\delta t}$:
$\mathcal{V}_{\delta t}$ =$=\Big{\{}\nu_{i}^{\delta t}\Big{\}}_{i\in\mathcal{I}}$ (45)
$\nu_{i}^{\delta t}$ $=var\left(\Delta x_{i}^{[1]},\dots,\Delta x_{i}^{[T-\delta t]}\right)$ (46)
where $var(\dots)$ represents the variance of a set of scalar variables.
Given a kernel function $K_{\alpha}(t_{1},t_{2})$, parameterized by the parameter $\alpha$, we can optimize the kernel parameter with a set of the variance datasets $\mathcal{V}_{\delta t}$ across different time step differences $\delta t$, through the following least-squares optimization problem:
$\alpha^{*}$ $=\operatorname*{arg\,min}_{\alpha}\sum_{\delta t}\left(c_{\alpha}^{\delta t}-\overline{\sum}_{i\in\mathcal{I}}\nu_{i}^{\delta t}\right)^{2}$ (47)
$c_{\alpha}^{\delta t}$ $=K_{\alpha}(0,0){-}K_{\alpha}(0,\delta t)K_{\alpha}(0,0)^{-1}K_{\alpha}(0,\delta t)$ (48)
We anecdotally demonstrate the sufficiency of our GP learning method by characterizing GP kernels from a human trajectory dataset [71] and a mouse behavior dataset [72]. The results are shown in Figure 4, where our method sufficiently distinguishes the difference in crowd navigation patterns between humans and mice.
V-C Condition GP on prior knowledge
The use of GP as the nominal mixed-strategy has one extra advantage: it allows us to condition uncertainty at specific time steps. Assuming we are interested in evaluating the nominal mixed-strategy at a set of time points $\mathbf{t}=[t_{1},\dots,t_{n}]$, we can transform the infinite-dimensional GP to a finite-dimensional multi-variate Gaussian distribution:
$\mathbf{s}$ $=[s(t_{1})\dots s(t_{n})]$
$\sim\mathcal{N}\left(\underbrace{[m(t_{1})\dots m(t_{n}))]}_{\mathbf{m}},\underbrace{\begin{bmatrix}K(t_{1},t_{1})\dots K(t_{1},t_{n})\\ \vdots\quad\ddots\quad\vdots\\ K(t_{n},t_{1})\dots K(t_{n},t_{n})\end{bmatrix}}_{\mathbf{\Sigma}}\right)$ (52)
However, following the stationary kernel assumption, the uncertainty of each time point in the above finite-dimensional Gaussian distribution (52) are all identical:
$\mathbf{K}_{i,i}=\mathbf{K}_{j,j},\quad\forall i,j\in[1,\dots,N]$ (53)
This property is suboptimal in practice, because we may have prior knowledge about the agent’s state uncertainty at certain time points. For example, the state uncertainty at the current time point should be smaller than the state uncertainty $5$ seconds later, or the state uncertainty when being close to a potential goal location should be smaller than the state uncertainty far from the goal location. Gaussian processes provide a natural solution to incorporate such information into uncertainty quantification. Given a set of time points $\bar{\mathbf{t}}=[\bar{t}_{1},\dots,\bar{t}_{m}]$ and the variance of the state uncertainty at these time points $\sigma=[\sigma_{1},\dots,\sigma_{n}]$, the new finite dimensional Gaussian distribution after incorporating the prior uncertainty knowledge is:
$\mathbf{s}$ $\sim\mathcal{N}\left(\mathbf{m},\mathbf{\Sigma}^{*}\right)$ (54)
$\mathbf{\Sigma}^{*}$ $=\mathbf{\Sigma}-\mathbf{\Sigma}^{\prime}(\overline{\mathbf{\Sigma}}+diag(\sigma))^{-1}\mathbf{\Sigma}^{\prime\top}$ (55)
where
$\mathbf{\Sigma}^{\prime}$ $=\begin{bmatrix}K(t_{1},\bar{t}_{1})&\dots&K(t_{1},\bar{t}_{m})\\ \vdots&\ddots&\vdots\\ K(t_{n},\bar{t}_{1})&\dots&K(t_{n},\bar{t}_{m})\end{bmatrix}$ (59)
$\overline{\mathbf{\Sigma}}$ $=\begin{bmatrix}K(\bar{t}_{1},\bar{t}_{1})&\dots&K(\bar{t}_{1},\bar{t}_{n})\\ \vdots&\ddots&\vdots\\ K(\bar{t}_{n},\bar{t}_{1})&\dots&K(\bar{t}_{n},\bar{t}_{n})\end{bmatrix}$ (63)
$diag(\sigma)$ $=\begin{bmatrix}\sigma_{1}&&\\ &\ddots&\\ &&\sigma_{n}\end{bmatrix}$ (67)
In Figure 3(a,b) and Figure 4, we condition the GP prior at the beginning and the end of the planning horizon with small uncertainty.
Remark 7.
The formula for conditioning GP covariance (55) is identical to the GP posterior in GP regression [73]. However, we are not conducting GP regression here since we assume the GP mean is given as a constant velocity model. Here we use the covariance formula in GP regression to characterize the uncertainty of the nominal mixed-strategies.
VI Sampling from mixed-strategy Nash equilibrium
Algorithm 2 describes how to find a Nash equilibrium for mixed-strategy crowd navigation, and Section V describes how to construct nominal mixed-strategy for each agent. However, real-time computation of Algorithm 2 is challenging, given the hardness of evaluating the high-dimensional integrals. Here we introduce sampling-based approximations of mixed-strategy Nash equilibrium, for the purpose of real-time path planning. The key insight for the real-time sampling approximation is to leverage the connection between Algorithm 2 and Bayesian updating and Monte-Carlo integration.
VI-A Sample-based mixed-strategy representation
Instead of maintaining mixed-strategies as continuous probability density functions, we now represent the mixed-strategy as a set of samples. We denote the sample set for each agent’s mixed-strategy as:
$\mathbf{p}_{i}=\{\mathbf{s}_{i,j}\}_{j=1,\dots,M}$ (68)
where $i$ is the agent index and $j$ is the sample index (we assume each agent has $M$ samples).
With the samples, we can asymptotically approximate the conditional likelihood (8) through Monte-Carlo integration:
$z(s|p_{/i})$ $=\exp\Big{(}{-}\mathbb{E}_{p_{/i}}[r](s)\Big{)}$ (69)
$\approx z(s|\mathbf{p}_{/i})$ (70)
$=\exp\left[{-}\overline{\sum_{i\in\mathcal{I}/i}}\left(\frac{1}{M}\sum_{j=1}^{M}r(s,\mathbf{s}_{i,j})\right)\right]$ (71)
Based on Bayes’ rule, for any sample $\mathcal{s}$,
VI-B Sample-based Nash equilibrium approximation
The approximated conditional likelihood (71) serves as the ratio between the posterior and prior distribution:
$\frac{p_{i}(s)}{p^{\prime}_{i}(s)}\approx\eta\cdot\exp\left[{-}\overline{\sum_{i\in\mathcal{I}/i}}\left(\frac{1}{M}\sum_{j=1}^{M}r(s,\mathbf{s}_{i,j})\right)\right]$ (72)
Furthermore, the closed-form asymptotic approximation (71) can be also considered as a ratio between target distribution (posterior) and proposal distribution (prior). Assuming samples from the prior distribution are accessible, we can apply tools from statistical simulation to approximate the propagation of mixed-strategies in Algorithm 2 with sample-based representation. Below we introduce two sampling schemes, which approximate Algorithm 2 through rejection sampling [74] and importance sampling [75], respectively.
Rejection sampling
The outline of the rejection sampling scheme is as follows:
  1. [nolistsep]
  2. 1.
    Draw sample $\mathbf{s}$ from the proposal distribution $p^{\prime}$;
  3. 2.
    Compute the ratio of the likelihood of measuring the sample in target distribution $p$ and proposal distribution $w=\frac{p(\mathbf{s})}{p^{\prime}(\mathbf{s})}$, as described in (72);
  4. 3.
    Uniformly draw a scalar sample $u$ between 00 and $1$;
  5. 4.
    If $\gamma u<w$, accept the sample $\mathbf{s}$, otherwise repeat the process. Here the user-specified parameter $\gamma>1$ serves as a constant, finite bound on the likelihood ratio $\frac{p(\mathbf{s})}{p^{\prime}(\mathbf{s})}$;
  6. 5.
    If trajectory prediction is required, compute the mean of the converged samples as the trajectory prediction.
Following this outline, the application of rejection sampling to Algorithm 2 is straightforward, since the likelihood ratio between the proposal distribution (nominal mixed-strategy) and target distribution (Bayesian posterior mixed-strategy) is already given in (71). A detailed pseudocode for the rejection sampling-based BRNE algorithm can be found in the appendix.
Importance sampling
Different from rejection sampling, importance sampling does not generate new samples. Instead, it assigns weights to samples from the proposal distribution (nominal mixed-strategy) to represent the target distribution (Bayesian posterior mixed-strategy). Denoting the weight for sample $\mathbf{s}_{i,j}$ as $\mathbf{w}_{i,j}$, the approximated conditional likelihood formula (71), with weighted samples, can be reformulated as:
$z(s|\mathbf{p}_{/i})$ $=\exp\left[{-}\overline{\sum_{i\in\mathcal{I}/i}}\left(\frac{1}{M}\sum_{j=1}^{M}\mathbf{w}_{i,j}\cdot r(s,\mathbf{s}_{i,j})\right)\right]$ (73)
We now give the outline of the importance sampling variant below:
  1. [nolistsep]
  2. 1.
    Draw sample $\mathbf{s}$ from the proposal distribution $p^{\prime}$. Initialize the weight for each sample as $1$;
  3. 2.
    Update sample weight with (73);
  4. 3.
    Once all sample weights are updated, normalize weights such that the averaged weight remains $1$;
  5. 4.
    If trajectory prediction is required, compute the weighted mean of the converged samples as the trajectory prediction.
A detailed pseudocode for the importance sampling-based BRNE algorithm can be found in the appendix.
Comparison of sampling schemes
Compared to important sampling, rejection sampling better approximates the posterior belief, since it re-generates new samples during iterations. However, the computation of rejection sampling is less efficient compared to importance sampling, due to the cost of re-generating samples. In practice, we use rejection sampling for explicit examination of the posterior beliefs (e.g., Figure 3(a,b) are generated in this way) and importance sampling for real-time planning of the robot (e.g., all the navigation experiments in this paper are conducted this way).
VI-C Sampling-based crowd navigation
We now describe the overall pipeline for robot crowd navigation supported by mixed-strategy Nash equilibrium:
  1. [nolistsep]
  2. 1.
    The robot perception stack returns the robot’s current position and surrounding pedestrians’ current positions and velocities;
  3. 2.
    Given a planning horizon, the robot computes the constant velocity model and combines it with a pre-trained GP kernel to generate nominal mixed-strategies for the surrounding humans. At the same time, the robot uses a meta-planner to plan a collision-free path within the environment (without humans) and combines it with the pre-trained GP kernel as the nominal mixed-strategy for the robot;
  4. 3.
    The robot runs sampling-based approximation of Algorithm 2, which returns a set of samples from the mixed-strategy Nash equilibrium among the robot and the surrounding humans;
  5. 4.
    The robot computes optimal control signals that follows the mean of the Nash equilibrium mixed-strategy, which is represented as a set of sample trajectories;
  6. 5.
    The robot executes the optimal control signals until the next replanning cycle, then repeat the above steps.
The optimal control signals from the Nash equilibrium mixed-strategy can be done through trajectory optimization or optimal control methods. We name our crowd navigation method Bayes’ rule Nash equilibrium (BRNE) navigation.
One of the most significant differences between our method and the existing social navigation methods is our method serves as a “correction” scheme on top of existing navigation methods, by taking the planned trajectories from these navigation methods as the nominal strategy for the robot. This design choice makes our method more flexible and easier to integrate into existing systems, such as the Robot Operating System (ROS) navigation stack [76, 77].
Figure 5: Examples of the joint navigation strategies (8 agents) from different methods, in the multi-agent navigation experiments.
TABLE I: Safety metrics of multi-agent navigation experiments.
#Agents
BRNE (RS)
BRNE (IS)
ORCA
CADRL
4 5 6 7 8
$1\%$ $1\%$ $2\%$ $4\%$ $4\%$
$2\%$ $3\%$ $4\%$ $5\%$ $7\%$
$\bf{0\%}$ $\bf{0\%}$ $\bf{0\%}$ $\bf{0\%}$ $\bf{0\%}$
$100\%$ $100\%$ $100\%$ $100\%$ $100\%$
[…]
4 5 6 7 8
$1.20{\pm}0.18$ $\bf{1.08{\pm}0.15}$ $\bf{1.00{\pm}0.13}$ $\bf{0.95{\pm}0.15}$ $\bf{0.93{\pm}0.14}$
$\bf{1.24{\pm}0.23}$ $1.07{\pm}0.18$ $0.96{\pm}0.18$ $0.87{\pm}0.15$ $0.76{\pm}0.13$
$0.6{\pm}0.0$ $0.6{\pm}0.0$ $0.6{\pm}0.0$ $0.6{\pm}0.0$ $0.6{\pm}0.0$
$0.17{\pm}0.15$ $0.08{\pm}0.07$ $0.05{\pm}0.04$ $0.04{\pm}0.03$ $0.04{\pm}0.02$
(b) Safety distance (m)—larger is better
TABLE II: Path efficiency of multi-agent navigation experiments.
#Agents 4 5 6 7 8
BRNE (RS) $6.44{\pm}0.06$ $\bf{6.61{\pm}0.09}$ $\bf{6.76{\pm}0.08}$ $\bf{6.84{\pm}0.08}$ $\bf{7.00{\pm}0.07}$
BRNE (IS) $6.90{\pm}0.32$ $7.06{\pm}0.34$ $7.23{\pm}0.33$ $7.36{\pm}0.32$ $7.36{\pm}0.31$
ORCA $\bf{6.38{\pm}0.35}$ $6.98{\pm}1.10$ $7.45{\pm}1.46$ $7.22{\pm}1.38$ $7.14{\pm}1.42$
CADRL $8.96{\pm}1.12$ $9.02{\pm}1.18$ $8.98{\pm}0.71$ $9.26{\pm}1.10$ $9.03{\pm}0.53$
(a) Maximum path length (m)—shorter is better
VI-D Algorithm complexity
We analyze the computational time complexity of a single iteration within the sampling-based algorithm. Given $T$ time steps as planning horizon, $N$ agents, and $M$ samples for each agent, the complexity of computing (72) for rejection-based sampling and (73) for importance sampling is $O(TMN)$. For each iteration of the Bayesian updating scheme, (72) and (73) need to be computed for all the agents, which leads to an overall complexity of $O(KMN^{2})$ for one iteration. On the other hand, the convergence rate of the iterative Bayesian updating scheme depends on the nominal mixed-strategies, which are unknown a priori. However, in Figure 6 we can see that the number of iterations required for the algorithm to converge does not depend on the number of players. Thus, our sampling-based crowd navigation in practice has a quadratic computation complexity of $O(KMN^{2})$ with respect to the number of agents. Note that, given $N$ agents, $O(N^{2})$ is the minimal complexity required for reasoning over all possible two-agent interaction pairs. The computation complexity of our algorithm is lower than the state-of-the-art dynamic game solver ALGAMES [60], which has a computation complexity of $O(MN^{3})$. Furthermore, the computation of sampling weights in (72) and (73) are highly parallelizable, which further reduces the computation time in practice.
VII Evaluation
VII-A Overview of evaluation
We evaluate our BRNE navigation model, as well as baseline methods, in two categories of tasks: (1) multi-agent navigation (Section VII-B); (2) crowd navigation (Section VII-C, Section VII-D). In the multi-agent navigation experiments, the tested method is asked to plan paths for all the agents, given a set of start locations and goal locations. Multi-agent navigation experiments represent the most ideal situation of crowd navigation, where all agents’ behaviors follow the assumption of the tested model. We evaluate multi-agent navigation performance as a lower bound of the algorithm’s ability to infer efficient and safe joint paths. However, for real-world navigation in human crowds, pedestrians do not necessarily follow the behavior assumed by the planning algorithm. Therefore, we conduct extensive crowd navigation experiments to test different algorithms’ performance in realistic crowd navigation scenarios. Furthermore, we conduct two kinds of crowd navigation experiments in different situations. The first crowd navigation evaluation (Section VII-C) places all the agents in a tight space where collision avoidance is necessary and simulates pedestrian behavior using a decentralized collision avoidance algorithm. The purpose of this evaluation is to stress test the robot’s ability to infer and adapt to cooperative but unknown joint navigation strategies from other agents. The second crowd navigation (Section VII-D) aims to benchmark the algorithm performance in realistic crowd navigation tasks. There, we conduct the evaluation in a state-of-the-art social navigation benchmark, which consists of real-world environmental geometry and real human navigation trajectories. The robot has to navigate in unstructured spaces while interacting with human pedestrians in non-controlled environments. For each experiment, we evaluate the algorithm performance from two categories of metrics: safety and navigation efficiency. Details of the specific metrics are provided in the corresponding subsections. We will release the code of our implementation.
VII-B Multi-agent navigation evaluation
Experiment design
For multi-agent navigation, we randomly place different numbers of agents on a circle with the radius being 3 meters. The goal of each agent is to reach the other side of the circle. This design means the most efficient path for each agent is to cross the center of the circle, which makes it necessary for all agents to compromise efficiency in exchange of safety. When randomizing the initial agent locations, we uniformly sample locations on the circle, but also make sure no two agents’ initial locations are within 0.6 meters of each other—we assume each agent is circular-shaped with a radius of 0.3 meters. We vary the number of agents from 4 to 8 and conduct 100 navigation trials for each number of agents. We specify the desired velocity for all agents and all baselines as $1.2m/s$. We test our method BRNE with both sampling schemes, we denote the rejection-sampling-based implementation as BRNE(RS) and the importance-sampling-based implementation as BRNE(IS).
Rationale of baselines
We select two baselines for comparison: (1) Optimal reciprocal collision avoidance (ORCA[8]; and (2) collision avoidance with deep reinforcement learning (CADRL[51]. We choose the selected baselines partially based on their widely accessible implementations. Furthermore, we choose ORCA as it is the de facto non-learning decentralized collision avoidance model for simulating crowds. We choose CADRL as it is the de facto learning-based crowd navigation model.
Rationale of metrics
For safety, we measure safety distance and collision rate. The safety distance is defined as the closest distance between any two agents during the whole navigation task. We consider a navigation trial involves collision if the safety distance of the trial is below 0.6 meters, twice the body radius of each agent. For efficiency, we measure the maximum path length among all agents in each trial.
Results
The evaluation results are presented in Table I(b) (safety) and Table II (navigation efficiency). Note that ORCA guarantees collision-free paths, thus it has $0\%$ collision rate across all experiments. On the other hand, our method BRNE generates joint navigation strategies with larger safety distance, better efficiency, and lower disparity. Even though BRNE strategies are not collision free, the collision rate is still minimal. Furthermore, as shown in Figure 5, ORCA is symmetry-sensitive, which leads to highly unnatural joint strategies when agents are unevenly distributed. Lastly, the learning-based baseline CADRL performs poorly across all metrics and exhibits highly non-smooth trajectories in Figure 5. Note that, the CADRL policy is trained using the toolbox provided in [52] with pedestrians simulated with ORCA. Thus, the trained CADRL policy is a collision avoidance policy in crowds. We test the CADRL policy in this multi-agent navigation benchmark to verify if the policy suffices as a coupled-prediction-and-planning method, instead of a dynamic object avoidance method. A detailed demonstration of the multi-agent navigation experiment results can be found in Extension 1.
Figure 6: Convergence of mixed-strategy across different numbers of agents in the multi-agent navigation experiments.
Figure 7: Plot of BRNE replanning frequencies with different numbers of agents and different numbers of samples per agent.
Computation time
We use the multi-agent navigation benchmark to also evaluate how the number of agents and the number of samples influence the computation time of BRNE. We fixed the planning horizon of BRNE at 50 time steps (5 seconds), then we vary the number of agents from 4 to 8 and the number of samples per agent from 100 to 500. The tested BRNE implementation uses Numba for parallelization and runs on a laptop with Intel Core i9-12900K Processor (20 threads). The results are shown in Figure 7. We can see that when the algorithm is accelerated through parallelization in practice, the computation complexity is close to linear with respect to the number of agents. In Table III, we further compare the computation speed of BRNE with two state-of-the-art dynamic game solvers, ALGAMES [60] and iLQGames [59]. We use 100 samples per agent for BRNE and use the results reported in the “intersection” experiment from [60]. Throughout this test, we use the importance-sampling-based BRNE implementation.
TABLE III: Comparison of computation time.
Algorithm 2 Agents 3 Agents 4 Agents
BRNE(Ours) $3.9{\pm}0.1$ ms $7.1{\pm}0.2$ ms $11.0{\pm}0.6$ ms
ALGAMES $50{\pm}11$ ms $116{\pm}22$ ms $509{\pm}33$ ms
iLQGames $752{\pm}168$ ms $362{\pm}93$ ms $1905{\pm}498$ ms
Convergence test
We use the multi-agent navigation benchmark to further test the convergence of the BRNE algorithm. In Figure 6, we show the convergence of BRNE with the number of agents varying from 4 to 8. We can see the algorithm consistently converges within 10 iterations and the convergence rates are similar across different numbers of agents. For better numerical accuracy, we use the rejection-sampling-based implementation for the convergence test.
VII-C Simulated crowd navigation evaluation
TABLE IV: Results of simulated crowd navigation experiments.
[…]
BRNE(Ours) $18\%$ $\bf{0.78{\pm}0.18}$ $8.29\pm 0.39$ $1.10\pm 0.09$
NCE $15\%$ $0.70\pm 0.08$ $\bf{8.12\pm 1.91}$ $1.15\pm 0.14$
ORCA $\bf{0\%}$ $0.6\pm 0.0$ $10.13\pm 0.70$ $\bf{1.02\pm 0.02}$
CADRL $96\%$ $0.41\pm 0.12$ $24.10\pm 0.00$ $1.29\pm 0.16$
Figure 8: Plots of time-to-goal metric vs. safety distance metric for all the methods. From left to right, the first three figures compare our method BRNE and other methods by overlapping data points from all 100 trials. The fourth figure overlaps the statistics (mean $\pm$ standard deviation) of all methods.
Figure 9: Frames from three example trials of our method BRNE in the simulated crowd navigation experiments. The blue circle with the letter “R” represents the robot and the orange circles with numbers represent the five simulated pedestrians. The robot’s task is to navigate to the goal indicated by the blue cross without running into pedestrians.
Experiment design
The geometrical layout of the simulated crowd navigation experiment is similar to the multi-agent navigation experiment, where we place agents across a circle with a radius of 3 meters and each agent’s task is to reach the other side of the circle. The difference is, in the crowd navigation experiment we only have control over one agent (the robot), while other agents are cooperative but make independent decisions. We use ORCA to simulate pedestrians since it is the most commonly used decentralized collision avoidance method. We test each navigation method with $5$ simulated pedestrians. The desired velocity for each agent is set as $1.2m/s$. The body radius of each agent is $0.3m$, thus the distance threshold for a collision is $0.6m$. We randomly initialize agent locations for 100 trials, with the same randomization scheme in the multi-agent navigation experiments. In this experiment, BRNE is implemented with importance sampling as a model predictive planner, updating its inference of mixed-strategy Nash equilibrium based on the latest observation of other agents’ positions and velocities.
TABLE V: Results of human dataset crowd navigation experiments.
Algorithm
Pedestrian collisions
Freezing behavior
Path length (m)
Time to goal (s)
BRNE(Ours) $\mathbf{1}$ $\mathbf{0}$ $16.56\pm 3.85$ $18.95\pm 4.81$
Meta-planner(Ours) $37$ $\mathbf{0}$ $\mathbf{15.42\pm 3.71}$ $17.66\pm 4.28$
SF $\mathbf{1}$ $1$ $17.25\pm 4.05$ $15.93\pm 4.17$
ORCA $15$ $1$ $17.66\pm 5.22$ $22.06\pm 7.78$
CADRL $40$ $1$ $15.70\pm 3.72$ $\mathbf{15.14\pm 4.21}$
Baseline $64$ $1$ $15.88\pm 3.57$ $16.08\pm 3.73$
Figure 10: Each row shows representative snapshots from a benchmark trial. The top row is from the UCY dataset and the bottom row is from the ETH dataset.
Figure 11: Snapshots from the only trial with collision. The collision is caused by the sudden appearance of a pedestrian right in front of the robot, leaving the robot no space and time to react.
Rationale of baselines
We select $3$ baselines to compare with: (1) ORCA; (2) CADRL; (3) Constrastive learning social navigation (NCE[54]. We choose ORCA and CADRL for the same reason as we choose them for multi-agent navigation evaluation. We choose NCE since it is reported as the state-of-the-art method in simulated crowd navigation benchmarks. We train NCE in the environment provided by [54].
Rationale of metrics
For safety, we measure the same safety distance and collision rate metrics. For navigation efficiency, we measure the time-to-goal and path length.
Results
In Table IV and Figure 8, we show the safety and navigation efficiency results of the simulated crowd navigation experiments. In Figure 9, we show frames of our method from the simulated tests. We can see that our method BRNE is competitive on both navigation safety and efficiency with the state-of-the-art reinforcement learning method NCE. Both NCE and CADRL are tested in the same simulation environment where it is trained, while our method BRNE requires no training. CADRL performs poorly in this benchmark, exhibiting high collision rate and low navigation efficiency. Furthermore, CADRL fails to reach the navigation goal and we have to terminate the trial after a certain amount of time, causing the zero standard deviation in the time-to-goal metric. The performance of CADRL could be due to the difference between the testing environment, which is provided by [54], and the training environment of CADRL. A detailed demonstration of the experimental results of our method BRNE in the simulated crowd navigation benchmark can be found in Extension 1.
VII-D crowd navigation in human datasets
Experiment design
We evaluate crowd navigation with prerecorded human pedestrian behaviors within the same geometrical spaces of recordings. The evaluation is conducted in SocNavBench [78], a state-of-the-art, high-fidelity social navigation benchmarking framework. SocNavBench contains human pedestrian data from two of the most commonly used pedestrian datasets ETH [79] and UCY [71]. A set of 33 curated episodes are extracted from the original datasets for benchmarking, which assemble a set of highly interactive test trials. The curated episodes contain an average of 44 pedestrians per episode. The distance between the goal location and the start location for the robot makes sure the navigation task can be finished within 25 seconds when the robot does not perform collision avoidance and navigates with the maximum permitted velocity of $1.2m/s$. An occupancy grid map of the real-world space from the original dataset is provided for each episode, the robot has access to the map a priori and can only navigate in the free space. Note that, even though the pedestrian agents in SocNavBench are non-reactive given the pre-recorded human behavior, the robot agent still assumes the pedestrian agents are real-world reactive agents. We refer readers to [78] for more details about the implementation of the benchmark, as well as the discussion on the choice of using pre-recorded human behavior. In this experiment, BRNE is implemented with importance sampling as a model predictive planner, updating its inference of mixed-strategy Nash equilibrium based on the latest observation of other agents’ positions and velocities.
Rationale of baselines
We compare our method BRNE with the four baselines reported in the original paper of SocNavBench [78], which are (1) the social force model (SF[7]; (2) ORCA; (3) CADRL; and (4) a pedestrian-unaware baseline with static obstacle avoidance capability (Baseline). Note that ORCA and CADRL do not have static obstacle avoidance capability, thus they use Baseline as a meta-planner. Similarly, we implement our meta-planner with static obstacle avoidance capability for BRNE based on a constant velocity model.
Rationale of metrics
We use the four main metrics based on the original metrics provided by the benchmark: (1) The total number of pedestrian collisions; (2) The total number of freezing behaviors, where the robot fails to reach the goal within 60 seconds or runs into environmental obstacles; (3) Path length; (4) Time to goal. Full results with complete benchmark metrics are provided in the appendix.
Results
Table V shows results of the human dataset crowd navigation experiments. Figure 10 shows representative snapshots of the benchmark tests. Both our method (BRNE) and the social force model (SF) significantly outperform the rest of the methods on navigation safety. Our method shares the minimal number of collisions with the social force model while being the only method with zero freezing behavior and exhibiting better path efficiency compared with the social force model. More importantly, compared to our meta-planner baseline Meta-planner(Ours), whose path is used by BRNE as the nominal strategy, BRNE significantly improves navigation safety, reducing the number of pedestrian collisions from 37 to 1, while sharing nearly identical navigation efficiency as the baseline. Further, as discussed in the next subsection, this one collision incident of our method is due to the pre-processing of the dataset instead of our method’s performance. These experimental results show that, when interacting with real human crowds in unstructured environments, our method is able to improve an existing navigation method’s safety to be near-perfect with minimal impact on navigation efficiency. This indicates that our method reaches human-level crowd navigation performance, on both safety and navigation efficiency, on top of a meta-planner, and also shows great potential for deploying our method on other existing navigation pipelines. A detailed demonstration of the human dataset benchmark can be found in Extension 1.
Fail case analysis
There is one trial where the robot collides with a pedestrian, here we provide an analysis of the cause, with the snapshots of the collision shown in Figure 11. As we can see, the sudden appearance of a pedestrian right in front of the robot, which is an artifact of the pre-processing of the dataset, leaves the robot with no space and no time to react. Therefore, this collision incident does not reflect our method’s navigation safety performance and can be avoided in practice. A detailed demonstration of the failure trial can be found in Extension 1.
Discussion on sim-to-real transfer
While we leave testing our method in real-world crowds for future work, due to the time and financial cost of conducting such experiments, we want to discuss the potential gap of sim-to-real transfer based on the human dataset benchmark results. First, since the robot is observing real-world human behavior in the benchmark, the benchmark results can at least verify which method will fail in the real world—a method that is dangerous or inefficient in the benchmark will likely remain dangerous or inefficient in the real world. Second, one main challenge of sim-to-real transfer in crowd navigation is the misalignment of human behavior assumptions. In our case, by inferring mixed-strategy Nash equilibrium for navigation, the robot assumes the humans are cooperative, which is not necessarily true in reality. However, the human dataset benchmark can serve as a preliminary verification of the influence of the misaligned human behavior assumption. Recall that, even though human agents are not reactive in the benchmark, the robot still believes it is interacting with real humans and is indeed observing real human behavior. The numerical efficiency of our method enables fast replanning during navigation, allowing the robot to constantly adjust its plan based on the latest observation, including observations that humans do not follow the assumed cooperative behavior. Furthermore, the robot makes navigation decisions based on mixed-strategy Nash equilibrium, it does not anticipate exact human behavior but instead a distribution of likely human behaviors, which already takes into account unexpected human reactions. Thus, the fast replanning capability and mixed-strategy-based modeling make the performance of our method in the human dataset benchmark more transferable in the real world. Lastly, our method is essentially built on top of an existing meta-planner. The modular design of our algorithm enhances the deployment flexibility, which lowers the gap of sim-to-real transfer. For example, our method can be integrated into existing robot navigation frameworks such as the navigation stack from the Robot Operating System. This is particularly important for crowd navigation in unstructured real-world environments, which most existing crowd navigation methods and all game-theoretic planning methods do not address. Lastly, even though out of the scope of this paper, we want to point out one important aspect of the sim-to-transfer that is not addressed in the benchmark: the perception aspect of crowd navigation. The benchmark assumes near-perfect robot perception of pedestrian states, but robust perception for both robot localization and pedestrian tracking remains an open challenge and could influence the real-world deployment of our method.
VIII Conclusion
In this work, we propose a computation-efficient mixed-strategy Nash equilibrium model for crowd navigation. Mixed-strategy Nash equilibrium provides a high-level cooperation model for the robot to plan actions that leverage human cooperation for collision avoidance while maintaining uncertainty in human behavior. Despite the general hardness of computing mixed-strategy Nash equilibrium, we achieve real-time inference speed on a laptop CPU by establishing the formal connection between mixed-strategy Nash equilibrium and a simple iterative Bayesian belief updating scheme. We further enhance the proposed Bayesian updating scheme by extracting temporal Gaussian processes from offline trajectory datasets as part of the Nash equilibrium model, using existing datasets to bring the model closer to actual human behavior. The proposed model can be incorporated into existing navigation frameworks to navigate alongside humans in unstructured environments with static obstacles. Our experiment results show that our model significantly improves the safety of a human-unaware planner without compromising navigation efficiency. Compared to other crowd navigation algorithms, our model consistently outperforms them in both safety and navigation efficiency. Our method also demonstrates human-level crowd navigation performance on top of a meta-planner when interacting with real human crowds in unstructured environments.
Beyond the technical contributions, our work also provides valuable insights into how game theory models can be applied to robotics applications. Most, if not all, existing game-theoretic planning algorithms are designed with a top-down approach: the planner aims to solve a generalized game, with the player objectives being specified later for different applications. While this approach can be applied to arbitrary games, the limitation of the top-down approach is the hardness of solving a generalized game, which is often intractable or too expensive to compute. In this work, we take a bottom-up approach to applying game theory models. We use mixed-strategy Nash equilibrium as a high-level principle to design a specific behavioral model for crowd navigation, with the purpose of supporting real-time decision-making. This bottom-up approach allows us to utilize the analytical power of mixed-strategy Nash equilibrium while maintaining a sufficient computation load for real-time robot navigation. We consider our bottom-up approach makes an orthogonal contribution, with respect to existing top-down approaches, to the principles of designing game-theoretic planners.
There are several limitations of the proposed crowd navigation framework. First, our framework does not support extra constraints on decision-making, such as guarantees on static obstacle avoidance. Even though we have zero static obstacle collision in the human dataset benchmark, this is achieved by using a collision-free meta-planner as the nominal strategy. Our method does not guarantee the robot from a collision with environmental obstacles. Second, the practical performance of our crowd navigation framework is limited by the sampling scheme. We proposed two sampling strategies for approximating the mixed-strategy Nash equilibrium. The rejection sampling-based strategy generates a more accurate posterior but is too slow for real-time decision-making. The importance-sampling-based strategy supports real-time posterior estimation, but its performance is limited by the support of initial samples and suffers from common issues faced by particle filters, such as sample degeneracy. Third, throughout the experiments, we assume the localization for the robot and perception of pedestrians are given, while perception in real human crowds is still an open challenge. Our ongoing work involves testing the algorithm on the hardware in the real world, and investigating how sensor noise would affect the algorithm’s performance.
Mixed-strategy Nash equilibrium paves the way for future works to achieve truly adaptive crowd navigation in varying environments, as it provides extra information enhancing the robot’s adaptivity during the run-time. We want to point out an exciting future direction stemming from our model, which is adaptive or contingent control synthesis based on mixed-strategy Nash equilibrium. In this work, we simply synthesize robot control signals to follow the mean of the converged mixed-strategy of the robot. This approach does not fully utilize the information encoded in mixed-strategy Nash equilibrium. For example, mixed-strategy predictions can serve as a prior distribution when existing information is insufficient for precise behavior prediction. Based on this prior, the robot could quickly refine its prediction or perform contingency planning with new online measurements.
Funding
This material is supported by the Honda Research Institute Grant HRI-001479. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the aforementioned institutions.
Declaration of conflicting interests
The author(s) declared no potential conflicts of interest with respect to the research, authorship, and/or publication of this article.
Appendix A: Proofs
Theorem 2 and Theorem 3 can be used to prove the convergence of Algorithm 1 and Algorithm 2. Again, we start with two-agent social navigation and extend the result to multi-agent crowd navigation.
Theorem 5.
The sequence of $\{(p_{i}^{[k]},p_{j}^{[k]})\}$ from the iterations of Algorithm 1 is a convergent sequence.
Proof.
We prove the convergence of the sequence through the monotone convergence theorem, where we show that the sequence monotonically decreases on a function with finite lower bound. We start with the monotone decrease result.
Based on the global optimality result in Theorem 2, for each iteration in Algorithm 1, after computing Agent $i$’s posterior belief $p_{i}^{[k+1]}$, we have the following inequality:
$\mathbb{E}_{p_{i}^{[k+1]},p_{j}^{[k]}}[r]{+}D(p_{i}^{[k+1]}\|p^{\prime}_{i})\leq\mathbb{E}_{p_{i}^{[k]},p_{j}^{[k]}}[r]{+}D(p_{i}^{[k]}\|p^{\prime}_{i})$ (74)
Adding the term $D(p_{j}^{[k]}\|p^{\prime}_{j})$ to both sides gives us:
$\mathbb{E}_{p_{i}^{[k+1]},p_{j}^{[k]}}[r]+D(p_{i}^{[k+1]}\|p^{\prime}_{i})+D(p_{j}^{[k]}\|p^{\prime}_{j})$
[…]
Then, given $p_{i}^{[k+1]}$, updating $p_{j}^{[k]}$ gives us the following inequality:
$\mathbb{E}_{p_{i}^{[k+1]},p_{j}^{[k+1]}}[r]{+}D(p_{j}^{[k+1]}\|p^{\prime}_{j})\leq\mathbb{E}_{p_{i}^{[k+1]},p_{j}^{[k]}}[r]{+}D(p_{j}^{[k]}\|p^{\prime}_{j})$ (76)
Adding $D(p_{i}^{[k+1]}\|p^{\prime}_{i})$ to both sides gives us:
$\mathbb{E}_{p_{i}^{[k+1]},p_{j}^{[k+1]}}[r]+D(p_{i}^{[k+1]}\|p^{\prime}_{i})+D(p_{j}^{[k+1]}\|p^{\prime}_{j})$
[…]
By applying the chain rule to inequality (75) and (77), we have:
$\mathbb{E}_{p_{i}^{[k+1]},p_{j}^{[k+1]}}[r]+D(p_{i}^{[k+1]}\|p^{\prime}_{i})+D(p_{j}^{[k+1]}\|p^{\prime}_{j})$
[…]
The above inequality means the sequence $\{(p_{i}^{[k]},p_{j}^{[k]})\}$ from Algorithm 1 monotonically decrease the function:
$R(p_{i}^{[k]},p_{j}^{[k]})+D(p_{i}^{[k]}\|p^{\prime}_{i})+D(p_{j}^{[k]}\|p^{\prime}_{j})$ (79)
Based on Assumption 1 and the non-negativity of KL-divergence, this function (79) has a finite lower-bound. Thus, based on the monotone convergence theorem, the sequence $\{(p_{i}^{[k]},p_{j}^{[k]})\}$ is convergent. ∎∎
Theorem 5 can be extended to the case of multi-agent crowd navigation in Algorithm 2.
Theorem 6.
The sequence of $\{(p_{0}^{[k]},\dots,p_{N-1}^{[k]})\}$ from the iterations of Algorithm 2 is a convergent sequence.
Proof.
At iteration $k$, for an arbitrary agent with index $a\in\mathcal{I}$, based on the global optimality result in Theorem 3, we have the following inequality:
${\sum_{i=1}^{a-1}}\mathbb{E}_{p_{a}^{[k+1]},p_{i}^{[k+1]}}[r]{+}{\sum_{i=a+1}^{N}}\mathbb{E}_{p_{a}^{[k+1]},p_{i}^{[k]}}[r]{+}D(p_{a}^{[k+1]}\|p^{\prime}_{a})$
[…]
By adding the following nested summation to both sides of the inequality:
$\sum_{i=1}^{a-1}\sum_{j=i+1}^{a-1}\mathbb{E}_{p_{i}^{[k+1]},p_{j}^{[k+1]}}[r]+\sum_{i=a+1}^{N}\sum_{j=a+1}^{N}\mathbb{E}_{p_{i}^{[k]},p_{j}^{[k]}}[r]$
$+\sum_{i=1}^{a-1}D(p_{i}^{[k+1]}\|p^{\prime}_{i})+\sum_{i=a+1}^{N}D(p_{i}^{[k]}\|p^{\prime}_{i})$ (81)
we have the left-hand side as:
${\sum_{i=1}^{a-1}}\mathbb{E}_{p_{a}^{[k+1]},p_{i}^{[k+1]}}[r]{+}{\sum_{i=a+1}^{N}}\mathbb{E}_{p_{a}^{[k+1]},p_{i}^{[k]}}[r]{+}D(p_{a}^{[k+1]}\|p^{\prime}_{a})$
$+{\sum_{i=1}^{a-1}\sum_{j=i+1}^{a-1}}\mathbb{E}_{p_{i}^{[k+1]},p_{j}^{[k+1]}}[r]{+}{\sum_{i=a+1}^{N}\sum_{j=a+1}^{N}}\mathbb{E}_{p_{i}^{[k]},p_{j}^{[k]}}[r]$
$+\sum_{i=1}^{a-1}D(p_{i}^{[k+1]}\|p^{\prime}_{i})+\sum_{i=a+1}^{N}D(p_{i}^{[k]}\|p^{\prime}_{i})$ (82)
=$\quad$= $\sum_{i=1}^{a-1}\sum_{j=i+1}^{a}\mathbb{E}_{p_{i}^{[k+1]},p_{j}^{[k+1]}}[r]+{\sum_{i=a+1}^{N}}R(p_{a}^{[k+1]},p_{i}^{[k]})$
$+\sum_{i=a+1}^{N}\sum_{j=a+1}^{N}\mathbb{E}_{p_{i}^{[k]},p_{j}^{[k]}}[r]$
$+\sum_{i=1}^{a}D(p_{i}^{[k+1]}\|p^{\prime}_{i})+\sum_{i=a+1}^{N}D(p_{i}^{[k]}\|p^{\prime}_{i})$ (83)
and the right hand side as:
${\sum_{i=1}^{a-1}}\mathbb{E}_{p_{a}^{[k]},p_{i}^{[k+1]}}[r]{+}{\sum_{i=a+1}^{N}}\mathbb{E}_{p_{a}^{[k]},p_{i}^{[k]}}[r]{+}D(p_{a}^{[k]}\|p^{\prime}_{a})$
$+{\sum_{i=1}^{a-1}\sum_{j=i+1}^{a-1}}\mathbb{E}_{p_{i}^{[k+1]},p_{j}^{[k+1]}}[r]{+}{\sum_{i=a+1}^{N}\sum_{j=a+1}^{N}}\mathbb{E}_{p_{i}^{[k]},p_{j}^{[k]}}[r]$
$+\sum_{i=1}^{a-1}D(p_{i}^{[k+1]}\|p^{\prime}_{i})+\sum_{i=a+1}^{N}D(p_{i}^{[k]}\|p^{\prime}_{i})$ (84)
=$\quad$= ${\sum_{i=1}^{a-1}\sum_{j=i+1}^{a-1}}\mathbb{E}_{p_{i}^{[k+1]},p_{j}^{[k+1]}}[r]{+}{\sum_{i=1}^{a-1}}\mathbb{E}_{p_{a}^{[k]},p_{i}^{[k+1]}}[r]$
$+{\sum_{i=a+1}^{N}\sum_{j=a}^{N-1}}\mathbb{E}_{p_{i}^{[k]},p_{j}^{[k]}}[r]$
$+\sum_{i=1}^{a-1}D(p_{i}^{[k+1]}\|p^{\prime}_{i})+\sum_{j=i+1}^{N}D(p_{i}^{[k]}\|p^{\prime}_{i})$ (85)
Combining both sides gives us the following inequality:
$\sum_{i=1}^{a-1}\sum_{j=i+1}^{a}\mathbb{E}_{p_{i}^{[k+1]},p_{j}^{[k+1]}}[r]+{\sum_{i=a+1}^{N}}R(p_{a}^{[k+1]},p_{i}^{[k]})$
$+\sum_{i=a+1}^{N}\sum_{j=a+1}^{N}\mathbb{E}_{p_{i}^{[k]},p_{j}^{[k]}}[r]$
$+\sum_{i=1}^{a}D(p_{i}^{[k+1]}\|p^{\prime}_{i})+\sum_{i=a+1}^{N}D(p_{i}^{[k]}\|p^{\prime}_{i})$
[…]
$+{\sum_{i=a+1}^{N}\sum_{j=a}^{N-1}}\mathbb{E}_{p_{i}^{[k]},p_{j}^{[k]}}[r]+\sum_{i=1}^{N}$
$+\sum_{i=1}^{a-1}D(p_{i}^{[k+1]}\|p^{\prime}_{i})+\sum_{j=i+1}^{N}D(p_{i}^{[k]}\|p^{\prime}_{i})$ (86)
By iterating the agent index $a$ from 00 to $N-1$, based on the chain rule of inequality, we have the following chain of inequalities:
$\sum_{i=1}^{N}\sum_{j=i+1}^{N}\mathbb{E}_{p_{i}^{[k+1]},p_{j}^{[k+1]}}[r]+\sum_{i=1}^{N}D(p_{i}^{[k+1]}\|p^{\prime}_{i})$
[…]
$+{\sum_{i=a+1}^{N}\sum_{j=a}^{N-1}}\mathbb{E}_{p_{i}^{[k]},p_{j}^{[k]}}[r]+\sum_{i=1}^{N}$
$+\sum_{i=1}^{a-1}D(p_{i}^{[k+1]}\|p^{\prime}_{i})+\sum_{j=i+1}^{N}D(p_{i}^{[k]}\|p^{\prime}_{i})$ (87)
[…]
which means the sequence $\{(p_{0}^{[k]},\dots,p_{N-1}^{[k]})\}$ is monotonically decreases the function:
$\sum_{i=1}^{N}\sum_{j=i+1}^{N}\mathbb{E}_{p_{i}^{[k]},p_{j}^{[k]}}[r]+\sum_{i=1}^{N}D(p_{i}^{[k]}\|p^{\prime}_{i})$ (89)
Based on Assumption 1 and the non-negativity of KL-divergence, this function (89) is lower-bounded. Thus, based on the monotone convergence theorem, the sequence $\{(p_{1}^{[k]},\dots,p_{N}^{[k]})\}$ is convergent. ∎∎
VIII-A Proof for Theorem 1
Proof.
We prove this by contradiction. Denote the sequence converges to the point $\{(p_{1}^{*},\dots,p_{N}^{*})\}$, assume the convergent point is not a Nash equilibrium, then:
$J_{i}(p_{1}^{*},\dots,p_{i},\dots,p_{N}^{*})<J_{i}(p_{1}^{*},\dots,p_{i}^{*},\dots,p_{N}^{*})$ (90)
$\exists p_{i}\in\mathcal{P},i\in\mathcal{I}$
Based on the definition of the general-sum game player objective (12), the above inequality means $p_{i}^{*}$ is not the global minimum of the following function:
$\mathbb{E}_{p,p_{/i}}[r]+D(p\|p^{\prime}_{i})$ (91)
This contradicts Theorem 3. ∎∎
VIII-B Proof for Theorem 4
Proof.
Recall in the proof for Theorem 6, the sequence of $\{(p_{0}^{[k]},\dots,p_{N-1}^{[k]})\}$ from the iterations of Algorithm 2 monotonically decreases the function:
$\sum_{i=1}^{N}\sum_{j=i+1}^{N}\mathbb{E}_{p_{i}^{[k]},p_{j}^{[k]}}[r]+\sum_{i=1}^{N}D(p_{i}^{[k]}\|p^{\prime}_{i})$ (92)
Since $p_{i}^{[0]}=p^{\prime}_{i}$, since KL-divergence is zero between two identical distributions, we have the following inequality:
$\sum_{i=1}^{N}\sum_{j=i+1}^{N}R(p_{i}^{*},p_{j}^{*})+\sum_{i=1}^{N}D(p_{i}^{*}\|p^{\prime}_{i})$
[…]
$=$ $\sum_{i=1}^{N}\sum_{j=i+1}^{N}R(p^{\prime}_{i},p^{\prime}_{j})+0$ (94)
Appendix B: Complete human dataset benchmark results
Complete results from the human dataset benchmark are shown in Table VI, Table VII, and Table VIII. Definitions of the metrics can be found in [78].
TABLE VI: Meta statistics of human dataset crowd navigation experiments.
Algorithm
Overall
Success rate
Failure cases
(T/PC/EC)
Total pedestiran
collisions
Planning wall time
per episode (s)
BRNE(Ours) $32/33$ $(0/1/0)$ $1$ $20.39\pm 6.93$
Baseline(Ours) $10/33$ $(0/23/0)$ $37$ $9.07\pm 2.12$
SF $32/33$ $(1/0/0)$ $1$ $18.23\pm 7.41$
ORCA $24/33$ $(1/8/0)$ $15$ $48.84\pm 23.06$
CADRL $18/33$ $(0/14/1)$ $40$ $46.78\pm 21.38$
Baseline $9/33$ $(1/23/0)$ $64$ $51.12\pm 16.21$
TABLE VII: Path quality results of human dataset crowd navigation experiments.
Algorithm
Path length (m)
Path length ratio
Goal traversal ratio
Path irregularity
Path traversal
time (s)
BRNE(Ours) $16.56\pm 3.85$ $1.06\pm 0.05$ $0.04$ $1.58\pm 1.04$ $18.95\pm 4.81$
Baseline(Ours) $15.42\pm 3.71$ $0.99\pm 0.03$ $0.02$ $1.58\pm 1.05$ $17.66\pm 4.28$
SF $17.25\pm 4.05$ $1.15\pm 0.23$ $0.52$ $1.66\pm 0.95$ $15.93\pm 4.17$
ORCA $17.66\pm 5.22$ $1.17\pm 0.26$ $0.21$ $1.56\pm 1.01$ $22.06\pm 7.78$
CADRL $15.70\pm 3.72$ $1.04\pm 0.05$ $0.51$ $1.68\pm 1.04$ $15.14\pm 4.21$
Baseline $15.88\pm 3.57$ $1.05\pm 0.11$ $0.09$ $1.65\pm 1.02$ $16.08\pm 3.73$
TABLE VIII: Motion quality results of human dataset crowd navigation experiments.
Algorithm
Average speed ($m/s$)
(max=$1.2m/s$)
Average energy
expenditure (J)
Average
acceleration ($m/s^{2}$)
Average jerk ($m/s^{3}$)
BRNE(Ours) $0.88\pm 0.31$ $327.97\pm 79.02$ $3.90\pm 7.76$ $151.44\pm 257.73$
Baseline(Ours) $0.88\pm 0.31$ $305.05\pm 74.25$ $3.91\pm 7.84$ $155.04\pm 260.59$
SF $1.09\pm 0.27$ $398.33\pm 95.45$ $0.39\pm 1.43$ $2.70\pm 28.23$
ORCA $0.80\pm 0.27$ $315.03\pm 89.89$ $0.31\pm 1.26$ $6.28\pm 28.85$
CADRL $1.04\pm 0.36$ $367.58\pm 87.24$ $0.93\pm 2.91$ $31.68\pm 87.71$
Baseline $0.99\pm 0.42$ $370.88\pm 84.55$ $4.81\pm 9.07$ $180.86\pm 303.88$
Appendix C: Pseudocode
See next page.
Algorithm 3 BRNE navigation (rejection sampling)
1:procedure BRNE_Nav($\mathbf{p}^{\prime}_{1},\dots,\mathbf{p}^{\prime}_{N},\gamma$) $\triangleright$ $\gamma>1$.
2:     $k\leftarrow 0$triangleright $k$ is the negotiation step.
3:     for $i\in[1,N]$ do
4:         for $j\in[1,M]$ do
5:              $\mathbf{s}_{i,j}^{[k]}\leftarrow\mathbf{s}^{\prime}_{i,j}$
6:         end for
7:     end for
8:     while convergence criterion not met do
9:         for $i\in[1,N]$ do
10:              $\mathbf{p}^{[k]}_{/i}\leftarrow\left({\bigcup_{a=1}^{i-1}}\mathbf{p}^{[k]}_{a}\right){\cup}\left(\bigcup_{a=i+1}^{N}\mathbf{p}^{[k+1]}_{a}\right)$
11:              for $j\in[1,M]$ do $\triangleright$ rejection sampling
12:                  $\mathbf{s}_{new}\leftarrow$ Draw new sample from $p^{\prime}_{i}$
13:                  $\omega\leftarrow z(\mathbf{s}_{new}|\mathbf{p}_{/i}^{[k]})$
14:                  $u\leftarrow Uniform(0,1)$
15:                  while $\gamma\cdot u\geq\omega$ do
16:                       $\mathbf{s}_{new}\leftarrow$ Draw new sample from $p^{\prime}_{i}$
17:                       $\omega\leftarrow z(\mathbf{s}_{new}|\mathbf{p}_{/i}^{[k]})$
18:                       $u\leftarrow Uniform(0,1)$
19:                  end while
20:                  $\mathbf{s}_{i,j}^{[k]}\leftarrow\mathbf{s}_{new}$
21:              end for
22:         end for
23:         $k\leftarrow k+1$
24:     end while
25:     return $\mathbf{p}_{1}^{[k]},\dots,\mathbf{p}_{N}^{[k]}$
26:end procedure
Algorithm 4 BRNE navigation (importance sampling)
1:procedure BRNE_Nav($\mathbf{p}^{\prime}_{1},\dots,\mathbf{p}^{\prime}_{N}$)
2:     $k\leftarrow 0$triangleright $k$ is the negotiation step.
3:     for $i\in[1,N]$ do
4:         for $j\in[1,M]$ do
5:              $\mathbf{w}_{i,j}^{[k]}\leftarrow 1$triangleright initialize sample weights
6:         end for
7:     end for
8:     while convergence criterion not met do
9:         for $i\in[1,N]$ do
10:              for $j\in[1,M]$ do $\triangleright$ importance sampling
11:                  $\mathbf{w}{\leftarrow}{\frac{1}{M}}{\sum_{a=1}^{i-1}}{\sum_{b=1}^{M}}\mathbf{w}_{a,b}^{[k]}r(\mathbf{s}_{i,j}^{\prime},\mathbf{s}_{a,b}^{\prime})$
12:                  $\mathbf{w}{\leftarrow}\mathbf{w}{+}{\frac{1}{M}}{\sum_{a=i+1}^{N}}{\sum_{b=1}^{M}}{\mathbf{w}_{a,b}^{[k+1]}}{r(\mathbf{s}_{i,j}^{\prime},\mathbf{s}_{a,b}^{\prime})}$
13:                  $\mathbf{w}_{i,j}^{[k+1]}\leftarrow\frac{1}{N-1}\mathbf{w}$
14:              end for
15:              Normalize $\mathbf{w}_{i}^{[k+1]}=[\mathbf{w}_{i,1}^{[k+1]},\dots,\mathbf{w}_{i,M}^{[k+1]}]$
16:         end for
17:         $k\leftarrow k+1$
18:     end while
19:     return $\mathbf{w}_{1}^{[k]},\dots,\mathbf{w}_{N}^{[k]}$
20:end procedure